### v1

@@ Abstract
We present a dynamic hedging scheme for S&P 500 options, where rebalancing decisions are enhanced by integrating information about the implied volatility surface dynamics. The optimal hedging strategy is obtained through a deep policy gradient-type reinforcement learning algorithm, with a novel hybrid neural network architecture improving the training performance. The favorable inclusion of forward-looking information embedded in the volatility surface allows our procedure to outperform several conventional benchmarks such as practitioner and smiled-implied delta hedging procedures, both in simulation and backtesting experiments.
@@ 1   Introduction
This paper builds on existing literature by developing a data-driven approach that incorporates IV surface information to derive optimal hedging positions. Unlike traditional methods focusing on local conditions, we adopt a global perspective, minimizing the risk metric associated with the terminal hedging error in a multi-period horizon framework. We leverage developments in risk-aware reinforcement learning (RL) to find the sequence of hedge ratios optimizing the hedger’s total risk until the option expiry, a setup analogous to Buehler et al., (2019)’s deep hedging approach. The novelty of our work lies in integrating factors that influence IV surface dynamics into the state variable set for determining hedging positions. The hedging agent therefore relies on market expectations for the underlying return distributions over several temporal horizons, producing genuine optimal forward-looking multi-stage hedging decisions informed by IV surfaces. To conduct our numerical experiments, we use the JIVR model of François et al., (2023), which is a parsimonious, tractable and data-driven econometric model representing the joint dynamics of the underlying asset return and five interpretable factors driving the IV surface of the S&P 500. Notably, the model has been calibrated on a data period spanning more than 25 years of daily option data, reflecting market behavior in a wide array of scenarios, including several financial crises.
As an additional contribution, we propose a novel hybrid neural network architecture combining feedforward and long-short-term memory (LSTM) layers, which is shown the improve training performance over conventional architecture.
@@
Hedging procedures with risk-aware reinforcement learning procedures have received substantial attention from the literature recently, see for instance Halperin, (2019), Cao et al., (2020), Du et al., (2020), Carbonneau and Godin, (2021), Carbonneau, (2021), Horvath et al., (2021), Imaki et al., (2021), Lütkebohmert et al., (2022), Cao et al., (2023), Carbonneau and Godin, (2023), Marzban et al., (2023), Mikkilä and Kanniainen, (2023), Pickard and Lawryshyn, (2023), Raj et al., (2023), Wu and Jaimungal, (2023) and Neagu et al., (2024). Nevertheless, to the best of our knowledge, our paper is the first to conduct multi-stage RL-based hedging that incorporates IV surface information directly as state variables.
@@
Our RL approach is shown to substantially outperform delta-based hedging strategies acting as benchmarks. In particular, RL agents trained with asymmetric objective functions such as the conditional value-at-risk (CVaR) or the semi-mean squared-error (SMSE) offer superior tradeoffs between profitability and downside risk. The outperformance of RL agents is even more pronounced in the presence of transaction costs, as such agents manage to develop hedging strategies that remain efficient in terms of risk mitigation while generating lower turnover. A feature importance analysis highlights that the conditional variance of the underlying asset returns, the level of the IV surface and its slope all significantly influence hedging performance, irrespective of the risk metric employed within the objective function.
@@
The rest of the paper is organized as follows. Section 2 frames the hedging problem in terms of a deep reinforcement learning framework. Section 3 provides the components of the JIVR model. Section 4 presents the numerical results, assessments, and global feature importance analysis. [^1] The Python code to replicate the numerical experiments from this paper can be found at the following link: https://github.com/cpmendoza/DeepHedging_JIVR.git. Section 5 concludes.
@@ 2.1   The hedging optimization problem
The European option payoff $\Psi(S_{T})$ depends on the price of the underlying asset at maturity, denoted as $T$ trading days. The hedging strategy involves managing a self-financing portfolio composed of both the underlying asset and a risk-free asset, with daily rebalancing. The strategy is represented by the predictable process $\{(\phi_{t},\delta_{t})\}_{t=1}^{T}$, where $\phi_{t}$ is the cash held at time $t-1$ and carried forward to the next period, and $\delta_{t}$ denotes the number of shares of the risky asset $S$ held during the interval $(t-1,t]$. The time-$t$ portfolio value is
@@
where $r_{t}$ is the time-$t$ annualized continuously compounded risk-free rate, $q_{t}$ is the annualized underlying asset dividend yield, both on the interval $(t-1,t]$, and $\Delta=\frac{1}{252}$. To account for transaction costs the self-financing condition entails
@@
Note that $\xi_{T}^{\delta}$ is a loss variable, with profits being represented by $-\xi_{T}^{\delta}$. The problem is
@@ 2.2.1   Neural network architecture
We propose a nonstandard architecture which consists in a Recurrent Neural Network with a Feedforward Connection (RNN-FNN) that combines the traditional Long Short-Term Memory Network (LSTM) and the Feedforward Neural Network (FFNN) architectures. The inclusion of LSTM layers mitigates problems related to vanishing gradients. [^2] Vanishing gradients arise when the gradients of the penalty function become extremely small, leading to slow or halted training (details can be found in Goodfellow et al., (2016)). The RNN-FNN network is defined as a composition of LSTM cells $\{C_{l}\}_{l=1}^{L_{1}}$ and FFNN layers $\{\mathcal{L}_{j}\}_{j=1}^{L_{2}}$ under the following functional representation:
@@
[…]
@@
[…]
@@
[…]
@@ 2.2.2   Neural network optimization
where $\theta_{j}$ is the set of parameters obtained after iteration $j$, $\eta_{j}$ is the learning rate (step size) which determines the magnitude of change in parameters on each time step, [^3] This parameter can be either deterministic or adaptive, i.e., it may be adjusted during the training period. For more details on this, please refer to Goodfellow et al., (2016). $\nabla_{\theta}$ is the gradient operator with respect to $\theta$ and $\hat{\mathcal{O}}$ is the Monte Carlo estimate of the objective function (3) computed on a mini-batch. Additional details are provided in Appendix A.
@@ 3   Joint market dynamics
This section describes the market dynamics that represent the joint evolution of the S&P 500 index price and its associated Implied Volatility (IV) surface. Such model is used to construct the state space of the hedging problem.
@@ 3.1   Daily implied volatility surface representation
On any given day, the cross-section of option prices on the S&P 500 index is captured through the IV surface model introduced by François et al., (2022) which characterizes the entire surface parsimoniously with a linear combination of five factors. More precisely, the time-$t$ IV of an option with time-to-maturity $\tau_{t}=\frac{T-t}{252}$ years and moneyness $M_{t}=\frac{1}{\sqrt{\tau_{t}}}\mbox{log}\frac{S_{t}e^{(r_{t}-q_{t})\tau_{t}}}{K}$, where $K$ is the strike price, is
@@
where $\beta_{t}=(\beta_{t,1},\beta_{t,2},\beta_{t,3},\beta_{t,4},\beta_{t,5})$ stands for the IV factor coefficients at time $t$ and the functions $\{f_{i}\}_{i=1}^{5}$ represent the long-term at-the-money (ATM) level, the time-to-maturity slope, the moneyness slope, the smile attenuation and the smirk, respectively (see Appendix B.1 for their specification).
Following the same data processing and estimation procedure outlined in the aforementioned study, we extract the daily time series of the IV factor coefficients spanning from January 4, 1996, to December 31, 2020.444The sample comes from OptionMetrics database, to which the conventional data exclusion filters are applied.
@@ 3.2   Joint Implied Volatility and Return (JIVR)
The first building block of the JIVR model represents the daily underlying asset excess log-return, $R_{t+1}=\mbox{log}\left(\frac{S_{t+1}}{S_{t}}\right)-(r_{t}-q_{t})\Delta$. It integrates an NGARCH(1,1) process with Normal Inverse Gaussian (NIG) innovations to capture volatility clusters and reproduce fat tailed asymmetric returns, while borrowing information from the volatility surface to anchor the evolution of the conditional variance process $h_{t,R}$. The second building block includes a multivariate heteroskedastic autoregressive processes with non-Gaussian innovations for all implied volatility factor coefficients. The multivariate time series representation of the JIVR model is presented in detail in Appendix B.2.
@@ 4.1.1   Market simulator
The JIVR model is used in subsequent simulation experiments to generate paths of the variables pertaining to market dynamics $(S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$, which drive the hedging decisions. The initial conditions ($\left(\{\beta_{0,i}\}^{5}_{i=1},h_{0,R},\{h_{0,i}\}^{5}_{i=1}\right)$ are randomly chosen among values prevailing in our data sample extending between January 4, 1996 and December 31, 2020. This constitutes a wide variety of states of the economy. Following the determination of initial values, the simulation progresses through all time steps in two distinct phases. First, a sequence of NIG innovations $\{\epsilon_{t}\}_{t=1}^{T}$ is simulated using the Monte Carlo method considering the dependence structure of contemporaneous innovations. Secondly, equations of Section B.2 are used to obtain values for ($\left(\{R_{t},\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1}\right)$ for $t=1,\ldots,T$ based on the simulated innovations.
@@ 4.1.2   Market parameters for numerical experiments
The initial underlying asset value is normalized to $S_{0}=100$ for simplicity. The options being hedged are assumed to be European call options ($\Psi(S_{T})=\mbox{max}(S_{T}-K,0)$) with maturities $T\in\{21,63,126\}$ days for short-term, medium-term and long-term maturities, and strikes $K\in\{90,100,110\}$ for in-the-money (ITM), at-the-money (ATM) and out-of-the-money (OTM) options, respectively. [^5] Unreported numerical results of European put options exhibit a similar pattern due to the Put-Call parity formula. Various levels of proportional transaction cost are considered, namely $\kappa\in\{0\%,0.05\%,0.5\%,1\%\}$. The initial value of the hedging strategy, $V_{0}^{\delta}$, is set to the price of the option being hedged, which is provided by the prevailing implied volatility surface. Other parameters of the model are specified in Section B.2.

### v2

@@ Abstract
We present a dynamic hedging scheme for S&P 500 options, where rebalancing decisions are enhanced by integrating information about the implied volatility surface dynamics. The optimal hedging strategy is obtained through a deep policy gradient-type reinforcement learning algorithm. The favorable inclusion of forward-looking information embedded in the volatility surface allows our procedure to outperform several conventional benchmarks such as practitioner and smiled-implied delta hedging procedures, both in simulation and backtesting experiments. The outperformance is more pronounced in the presence of transaction costs.
@@ 1  Introduction
This paper builds on existing literature by developing an approach that incorporates IV surface information to derive optimal hedging positions. Unlike traditional methods focusing on local conditions, we adopt a global perspective, minimizing the risk metric associated with the terminal hedging error in a multi-period horizon framework. We leverage developments in risk-aware reinforcement learning (RL) to find the sequence of hedge ratios optimizing the hedger’s total risk until the option expiry, a setup analogous to Buehler et al., (2019)’s deep hedging approach. The novelty of our work lies in integrating factors that influence IV surface dynamics into the state variable set for determining hedging positions. The hedging agent therefore relies on market expectations for the underlying return distributions over several temporal horizons, producing genuine optimal forward-looking multi-stage hedging decisions informed by IV surfaces. To conduct our numerical experiments, we use the JIVR model of François et al., (2023), which is a parsimonious and tractable econometric model representing the joint dynamics of the underlying asset return and five interpretable factors driving the IV surface of the S&P 500. Notably, the model has been estimated on a data period spanning more than 25 years of daily option data, reflecting market behavior in a wide array of scenarios, including several financial crises.
@@
Hedging procedures with risk-aware reinforcement learning procedures have received substantial attention from the literature recently, see for instance the following non-exhaustive list: Halperin, (2019), Cao et al., (2020), Du et al., (2020), Carbonneau and Godin, (2021), Carbonneau, (2021), Giurca and Borovkova, (2021), Horvath et al., (2021), Imaki et al., (2021), Lütkebohmert et al., (2022), Xu and Dai, (2022), Cao et al., (2023), Carbonneau and Godin, (2023), Marzban et al., (2023), Mikkilä and Kanniainen, (2023), Pickard and Lawryshyn, (2023), Raj et al., (2023), Wu and Jaimungal, (2023) and Neagu et al., (2024). Nevertheless, to the best of our knowledge, our paper is the first to conduct multi-stage RL-based hedging that incorporates IV surface information directly as state variables.
This study focuses on hedging procedures using the underlying asset as the only hedging instrument. Hedging strategies that add other instruments such as options potentially improve the quality of the hedge (by managing other Greeks like the gamma or the vega) but they come with higher transaction costs. In addition, they require a precise timely execution of trades between the option and the underlying asset markets. As a matter of fact, delta hedging has been a long-lasting risk management problem that continues to draw the attention of researchers (Hull and White,, 2017). It remains relevant for several market participants including option market makers (Huh et al.,, 2015; Kokholm and Egebjerg,, 2024), although the extent of their delta hedging has been recently challenged (Hu et al.,, 2023), insurance companies (Augustyniak and Boudreault,, 2012; Sun and Platen,, 2019), structured products issuers (Entrop and Fischer,, 2020; Henderson et al.,, 2020), or proprietary trading firms (Ni et al.,, 2005; O’Donovan et al.,, 2023).
@@
Our RL approach is shown to substantially outperform delta-based hedging strategies acting as benchmarks. In particular, RL agents trained with asymmetric objective functions such as the conditional value-at-risk (CVaR) or the semi-mean squared-error (SMSE) offer superior tradeoffs between profitability and downside risk. The outperformance of RL agents is even more pronounced in the presence of transaction costs, as such agents manage to develop hedging strategies that remain efficient in terms of risk mitigation while generating lower turnover. A variable importance analysis highlights that the conditional variance of the underlying asset returns, the level of the IV surface and its slope all significantly influence hedging performance, irrespective of the risk metric employed within the objective function.
The out-of-sample backtesting study highglights that RL-based hedging strategies can outperform traditional benchmarks, particularly when the algorithm is given access to rich market information such as the implied volatility (IV) surface. In the presence of transaction costs, RL approaches demonstrate superior hedging performance. However, in the absence of transaction costs, the benefits of RL hedging are contingent on the quality of the informational inputs—without IV surface data, the RL agent fails to outperform delta hedging in terms of mean squared error. These findings highlight the potential of reinforcement learning as a flexible and powerful tool for risk management, provided that it is informed with relevant market signals.
@@
The rest of the paper is organized as follows. Section˜2 frames the hedging problem in terms of a deep reinforcement learning framework. Section˜3 provides the components of the JIVR model. Section˜4 presents the numerical results, assessments, and global feature importance analysis. [^1] The Python code to replicate the numerical experiments from this paper can be found at the following link: https://github.com/cpmendoza/DeepHedging_JIVR.git. Section˜5 presents the out-of-sample backtest study. Section˜6 concludes.
@@ 2.1  The hedging optimization problem
The European option payoff $\Psi(S_{T})$ depends on the price of the underlying asset at maturity, denoted as $T$ trading days. The hedging strategy involves managing a self-financing portfolio composed of both the underlying asset and a risk-free asset, with daily rebalancing. The strategy is represented by the predictable process $\{(\phi_{t},\delta_{t})\}_{t=1}^{T}$, where $\phi_{t}$ is the cash held at time t1t-1 and carried forward to the next period, and $\delta_{t}$ denotes the number of shares of the risky asset $S$ held during the interval (t1,t](t-1,t]. The time-$t$ portfolio value is
@@
where $r_{t}$ is the time-$t$ annualized continuously compounded risk-free rate, $q_{t}$ is the annualized underlying asset dividend yield, both on the interval (t1,t](t-1,t], and $\Delta=\frac{1}{252}$. To account for transaction costs the self-financing condition entails
@@
Note that $\xi_{T}^{\delta}$ is a loss variable, with profits being represented by $\xi_{T}^{\delta}$. The problem is
@@ 2.2.1 Neural network architecture
We propose an architecture reminiscent of these used in Koochali et al., (2019) and Fecamp et al., (2020), which consists in a Recurrent Neural Network with a Feedforward Connection (RNN-FNN) that combines the traditional Long Short-Term Memory Network (LSTM) and the Feedforward Neural Network (FFNN) architectures. The inclusion of LSTM layers mitigates problems related to vanishing gradients. [^2] Vanishing gradients arise when the gradients of the penalty function become extremely small, leading to slow or halted training (details can be found in Goodfellow et al., (2016)). The RNN-FNN network is defined as a composition of LSTM cells $\{C_{l}\}_{l=1}^{L_{1}}$ and FFNN layers $\{\mathcal{L}_{j}\}_{j=1}^{L_{2}}$ under the following functional representation:
@@
The LSTM cell $C_{l}$ is a mapping that transforms a vector $Z^{(C,\,l-1)}_{t}$ of dimension $d^{(C,\,l-1)}$ into a vector $Z^{(C,\,l)}_{t}$ of dimension $d^{(C,\,l)}$ based on the following equations, considering $Z^{(C,\,0)}_{t}=X_{t}$:
@@
where $\mbox{sigm}(\cdot)$ and $\mbox{tanh}(\cdot)$ are respectively the sigmoid and hyperbolic tangent functions applied element-wise and $\odot$ is the Hadamard product. Layer $\mathcal{L}_{j}$ represents a FFNN layer that maps the input vector $Z^{(\mathcal{L},\,j-1)}_{t}$ of dimension $d^{(\mathcal{L},\,j-1)}$ into a vector $Z^{(\mathcal{L},\,j)}_{t}$ of dimension $d^{(\mathcal{L},\,j)}$ by applying a linear transformation $T_{\mathcal{L}_{j}}(Z^{(\mathcal{L},\,j-1)}_{t})=W_{\mathcal{L}_{j}}Z^{(\mathcal{L},\,j-1)}_{t}+b_{\mathcal{L}_{j}}$ and, subsequently, an element-wise non-linear activation function $g_{\mathcal{L}_{j}}$, i.e., $\mathcal{L}_{j}(Z^{(\mathcal{L},\,j-1)}_{t})=(g_{\mathcal{L}_{j}}\circ T_{\mathcal{L}_{j}})(Z^{(\mathcal{L},\,j-1)}_{t})$ for $j\in\{1,...,L_{2},J\}$, considering $Z^{(\mathcal{L},\,0)}_{t}=Z^{(C,\,L_{1})}_{t}$.
@@
  • If $L_{1}\geq l\geq 1$: $W_{i}^{(l)}$, $W_{o}^{(l)}$, $W_{c}^{(l)}\in\mathbb{R}^{d^{(C,\,l)}\times d^{(C,\,l-1)}}$ and $b_{i}^{(l)}$, $b_{o}^{(l)}$, $b_{c}^{(l)}\in\mathbb{R}^{d^{(C,\,l)}\times 1}$ with $d^{(C,\,0)}$ defined as the original dimension of the network input.
  • If $L_{2}\geq j\geq 1$: $W_{\mathcal{L}_{j}}\in\mathbb{R}^{d^{(\mathcal{L},\,j)}\times d^{(\mathcal{L},\,j-1)}}$ and $b_{\mathcal{L}_{j}}\in\mathbb{R}^{d^{(\mathcal{L},\,j)}}$ with $d^{(\mathcal{L},\,0)}=d^{(C,\,L_{1})}$.
@@ 2.2.2 Neural network optimization
where $\theta_{j}$ is the set of parameters obtained after iteration $j$, $\eta_{j}$ is the learning rate (step size) which determines the magnitude of change in parameters on each time step, [^3] This parameter can be either deterministic or adaptive, i.e., it may be adjusted during the training period. For more details on this, please refer to Goodfellow et al., (2016). $\nabla_{\theta}$ is the gradient operator with respect to $\theta$ and $\hat{\mathcal{O}}$ is the Monte Carlo estimate of the objective function (3) computed on a mini-batch. Additional details are provided in Appendix B of the supplementary material.
@@ 3  Joint market dynamics
This section describes the JIVR model of François et al., (2023), which provides market dynamics that we use to represent the joint evolution of the S&P 500 index price and its associated Implied Volatility (IV) surface. Such model is used to construct the state space of the hedging problem. The JIVR model takes the form of an econometric model directly specifying the evolution of the shape of the volatility surface. [^4] The direct modeling of the volatility surface is aligned with the so-called ”instrumental approach” philosophy detailed in Rebonato, (2004), which exhibits more flexibility and treats volatility surfaces as objects possessing intrinsic characteristics on their own, not being fully characterized by the underlying asset. The model has been crafted through the careful analysis of option price data spanning multiple economic periods and crises, with factor loadings being designed to mimic the most important loadings from a principal component analysis on volatility surfaces. The advantages of the JIVR model are twofold. First, the empirical approach used for its construction, which relies on daily fitting to IV data, enables the accurate representation of all segments on the IV surface on a daily basis. Second, it is very tractable and does not rely on the filtering of latent factors driving volatility dynamics as in multiple competing models that rely on risk-neutral dynamics to construct volatility surfaces.
@@ 3.1  Daily implied volatility surface representation
On any given day, the cross-section of option prices on the S&P 500 index is captured through the IV surface model introduced by François et al., (2022) which characterizes the entire surface parsimoniously with a linear combination of five factors. More precisely, the time-$t$ IV of an option with time-to-maturity $\tau_{t}=\frac{T-t}{252}$ years and (scaled) moneyness $M_{t}=\frac{1}{\sqrt{\tau_{t}}}\mbox{log}\frac{S_{t}e^{(r_{t}-q_{t})\tau_{t}}}{K}$, where $K$ is the strike price, is
@@
where $\beta_{t}=(\beta_{t,1},\beta_{t,2},\beta_{t,3},\beta_{t,4},\beta_{t,5})$ stands for the IV factor coefficients at time $t$ and the functions $\{f_{i}\}_{i=1}^{5}$ represent the long-term at-the-money (ATM) level, the time-to-maturity slope, the moneyness slope, the smile attenuation and the smirk, respectively (see Appendix C.1 of the supplementary material for their specification). Following the same data processing and estimation procedure outlined in the aforementioned study, we extract the daily time series of the IV factor coefficients spanning from January 4, 1996, to December 31, 2020.555The sample comes from OptionMetrics database, to which the conventional data exclusion filters are applied.
@@ 3.2  Joint Implied Volatility and Return (JIVR)
The first building block of the JIVR model represents the daily underlying asset excess log-return, $R_{t+1}=\mbox{log}\left(\frac{S_{t+1}}{S_{t}}\right)-(r_{t}-q_{t})\Delta$. It integrates an NGARCH(1,1) process with Normal Inverse Gaussian (NIG) innovations to capture volatility clusters and reproduce fat tailed asymmetric returns, while borrowing information from the volatility surface to anchor the evolution of the conditional variance process $h_{t,R}$. The second building block includes a multivariate heteroskedastic autoregressive processes with non-Gaussian innovations for all implied volatility factor coefficients. The multivariate time series representation of the JIVR model is presented in detail in Appendix C.2 of the supplementary material.
@@ 4.1.1 Market simulator
The JIVR model is used in subsequent simulation experiments to generate paths of the variables pertaining to market dynamics $(S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$, which drive the hedging decisions. [^6] Coefficient $\beta_{t-1,2}$ is omitted from the state space due to its minor impact. The initial conditions $\left(\{\beta_{0,i}\}^{5}_{i=1},h_{0,R},\{h_{0,i}\}^{5}_{i=1}\right)$ are randomly chosen among values prevailing in our data sample extending between January 4, 1996 and December 31, 2020. This constitutes a wide variety of states of the economy. Following the determination of initial values, the simulation progresses through all time steps in two distinct phases. First, a sequence of NIG innovations $\{\epsilon_{t}\}_{t=1}^{T}$ is simulated using the Monte Carlo method considering the dependence structure of contemporaneous innovations. Second, equations of Section C.2 from the supplementary material are used to obtain values for $\left(R_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1}\right)$ for $t=1,\ldots,T$ based on the simulated innovations.
@@ 4.1.2 Market parameters for numerical experiments
The initial underlying asset value is normalized to $S_{0}=100$ for simplicity. The options being hedged are assumed to be European call options ($\Psi(S_{T})=\mbox{max}(S_{T}-K,0)$) with maturities $T\in\{21,63,126\}$ days for short-term, medium-term and long-term maturities, and strikes $K\in\{90,100,110\}$ for in-the-money (ITM), at-the-money (ATM) and out-of-the-money (OTM) options, respectively. [^7] Unreported numerical results of European put options exhibit a similar pattern due to the Put-Call parity formula. Various levels of proportional transaction cost are considered, namely $\kappa\in\{0\%,0.05\%,0.5\%,1\%\}$. The initial value of the hedging strategy, $V_{0}^{\delta}$, is set to the price of the option being hedged, which is provided by the prevailing implied volatility surface. The annualized continuously compounded risk-free rate and dividend yield are assumed to be constant and are estimated using data from the period 2016 to 2020, with values set to $r=2.66\%$ and $q=1.77\%$, respectively. Other parameters of the model are specified in Section C.2.